OpenStudy (anonymous):

Translate the sentence into an equation The difference of a number times 2 and 7 is 5. Use the variable w for the unknown number.

OpenStudy (anonymous):

This is kinda a bad question...It defiantly means (Garbage)=5 -w(2+7)=5 maybe

OpenStudy (anonymous):

I put 2w-7=5

OpenStudy (anonymous):

but im not sure

OpenStudy (anonymous):

Altanta's answer is better.

OpenStudy (anonymous):

Okay :) thanks.